Evaluate the integral of (e^((¼)x+3) + 3x⁴)dx


1
Expert's answer
2021-03-31T16:29:46-0400
"\int(e^{( (1/4)x+3)}+3x^4)dx=4e^{( (1/4)x+3)}+\dfrac{3}{5}x^5+C"